OpenStudy (anonymous):

How do I change this equation to slope-intercept form? Explanation needed, thank you!

OpenStudy (anonymous):

this??

OpenStudy (tkhunny):

Solve for y. That's it.

OpenStudy (anonymous):

2x + 3y = 18 I feel dumb, sorry I didn't include this.

OpenStudy (tkhunny):

Oh, in that case, solve for y. Go!

OpenStudy (anonymous):

What would X be when I solve it?

OpenStudy (tkhunny):

x is just x. Why does it need a value?

OpenStudy (anonymous):

Would y be 6?

OpenStudy (amistre64):

no, just let x be x, rearrange the terms so that y is all by itself: like y = mx + b

OpenStudy (tkhunny):

When 'solving for" it doesn't go away. You still have it when you are done. a + b + c = 2 Solve for a a = 2 - b - c See. Everyone is still there. Solve for b b = 2 - a - c Again, no one vanished and it was not required to assign any specific value. It is an abstraction. You have to think of the letters as representing SOME number, but it is not necessary to know what that number is. Just have a little faith that one day you will know if it is necessary to your mathematical salvation. :-) Note: I am joking a little, but this is real. It is an abstraction. It might be a different way of thinking. You do have to get used to the idea. It's okay if it is a little confusing at first. You'll get it.